# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

### v1

@@ Appendix A Tag Extraction
TABLE VI: Ablation on number of concepts $K$.
$\alpha_{1}$ $\alpha_{2}$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
@@
TABLE VII: Ablation on loss parameters $\alpha_{1}$ and $\alpha_{2}$.
@@
Video-LLaMA [38] Visual 8,049 3,500 27,11 27,12
VideoLLaMA2 [37] Visual 63,383 12,118 27.69 27.83
Llama2 [40] Textual 162,571 5,058 14.96 15.20
Llama3.1 [95] Textual 320,351 8,326 27.17 26.52
@@
TABLE VIII: Comparison of statistics of tags when using different foundation models on MSR-VTT-9k.
DiCoSA* [3] - IJCAI’23 47.2 73.5 83.0 2 12.9 203.7 44.4 72.6 81.2 2 14.7 198.2 41.2 71.3 81.3 2 15.9 193.8
DiCoSA* [3] QB IJCAI’23 48.0 74.6 84.3 2 12.9 206.9 44.7 73.3 81.3 2 14.3 199.3 43.7 73.2 81.7 2 16.8 198.6
DiCoSA* [3] DSL IJCAI’23 52.1 77.3 85.9 1 12.9 215.3 49.2 73.3 82.8 2 12.6 205.3 47.3 75.7 83.8 2 14.2 206.8
DiCoSA* [3] - IJCAI’23 14.0 28.5 36.1 30 296.7 78.6 16.7 35.9 47.1 12 61.7 99.7 11.9 29.2 40.3 18 117.9 81.4
DiCoSA* [3] QB IJCAI’23 14.1 28.4 36.4 30 313.6 78.9 18.8 39.5 49.0 11 60.8 107.3 12.9 31.5 42.3 16 108.7 86.7
DiCoSA* [3] DSL IJCAI’23 15.5 30.1 37.8 28 290.6 83.4 18.2 39.2 49.1 11 58.9 106.5 12.8 32.2 43.0 15 103.8 88.0
TABLE IX: Comparison with baseline trained by using same training parameters of MAC-VR. our reproduced results with same training parameters. IS: Inference Strategy. $BS$: Batch Size. $N_{v}$: Number of Frames. The highest $RSum$ in each block is marked in bold.
Appendix B Full Tables of Experiments
Tab. IX details the results presented in Fig. 4 and reports Recall at $L=1,5,10$ ($R@L$), median rank $MR$, mean rank ($MeanR$) and $RSum$ for all the datasets when using all the inference strategies. Similarly, . X reports the full table of Tab. II. Tabs. XIIIXIII are the full tables of the ablation s. VV in the main paper.
MSR-VTT-9k MSR-VTT-7k DiDeMo
CLIP4Clip [77] - Neurocomp.’22 44.5 71.4 81.6 2 15.3 197.5 42.1 71.9 81.4 2 16.2 195.4 43.4 70.2 80.6 2 17.5 194.2
CenterCLIP [78] - SIGIR’22 44.2 71.6 82.1 2 15.1 197.9 43.7 71.3 80.8 2 16.9 195.8 - - - - - -
X-Pool [26] - CVPR’22 46.9 72.8 82.2 2 14.3 201.9 43.9 72.5 82.3 2 14.6 198.7 - - - - - -
TS2-Net [13] - ECCV’22 47.0 74.5 83.8 2 13.0 205.3 - - - - - - 41.8 71.6 82.0 2 14.8 195.4
EMCL-Net [15] - NeurIPS’22 46.8 73.1 83.1 2 - 203.0 - - - - - - 46.8 74.3 83.1 2 12.3 204.2
EMCL-Net [15] - NeurIPS’22 - - - - - - 45.2 71.4 81.2 2 15.0 197.8 44.1 71.7 80.3 2 15.1 197.8
VoP [27] - CVPR’23 44.6 69.9 80.3 2 16.3 194.8 42.7 68.2 79.3 2 15.9 190.2 46.4 71.9 81.5 2 13.6 199.8
MuMUR [79] - Inf. Retr. J.’23 46.4 72.6 82.2 2 13.9 201.2 44.8 72.0 82.5 2 - 199.3 44.4 74.3 83.1 2 - 201.8
PiDRo [29] - ICCV’23 48.2 74.9 83.3 2 12.6 206.4 - - - - - - 48.6 75.9 84.4 2 11.8 208.9
HBI [80] - CVPR’23 48.6 74.6 83.4 2 12.0 206.6 - - - - - - 46.9 74.9 82.7 2 12.1 204.5
DiffusionRet [14] - ICCV’23 49.0 75.2 82.7 2 12.1 206.9 - - - - - - 46.7 74.7 82.7 2 14.3 204.1
Prompt Switch [81] - ICCV’23 47.8 73.9 82.2 - 14.4 203.9 - - - - - - - - - - - -
Cap4Video [48] - CVPR’23 49.3 74.3 83.8 2 12.0 207.4 - - - - - - 52.0 79.4 87.5 1 10.5 218.9
UCoFiA [82] - ICCV’23 49.4 72.1 83.5 2 12.9 205.0 - - - - - - 46.5 74.8 84.4 2 13.4 205.7
UCoFiA [82] - ICCV’23 - - - - - - 45.2 69.6 79.4 2 15.7 194.2 42.1 69.2 79.1 2 16.3 190.4
PAU [83] - NeurIPS’23 48.5 72.7 82.5 2 14.0 203.7 - - - - - - 48.6 76.0 84.5 2 12.9 209.1
TABLE [10] - AAAI’23 47.1 74.3 82.9 2 13.4 204.3 - - - - - - 47.9 74.0 82.1 2 14.3 204.0
UATVR [33] - ICCV’23 47.5 73.9 83.5 2 12.3 204.9 - - - - - - 43.1 71.8 82.3 2 15.1 197.2
TeachCLIP [30] - CVPR’24 46.8 74.3 - - - - - - - - - - 43.7 71.2 - - - -
MV-Adapter [31] - CVPR’24 46.2 73.2 82.7 - - 202.1 - - - - - - 44.3 72.1 80.5 - - 196.9
BiC-Net [84] - TOMM’24 39.4 75.5 86.7 2 - 201.6 32.8 68.2 82.4 3 - 183.4 - - - - - -
RAP [85] - ACL’24 44.8 71.4 81.5 - 14.4 197.7 - - - - - - 42.6 70.4 79.6 - 18.0 192.6
QB-Norm [17] QB CVPR’22 47.2 73.0 83.0 2 - 203.2 - - - - - - 43.3 71.4 80.8 2 - 195.5
DiCoSA [3] QB IJCAI’23 47.5 74.7 83.8 2 13.2 206.0 - - - - - - 45.7 74.6 83.5 2 11.7 203.8
DiffusionRet [14] QB ICCV’23 48.9 75.2 83.1 2 12.1 207.2 - - - - - - 48.9 75.5 83.3 2 14.1 207.7
EMCL-Net [15] DSL NeurIPS’22 51.6 78.1 85.3 1 - 215.0 - - - - - - - - - - - -
EMCL-Net [15] DSL NeurIPS’22 - - - - - - - - - - - - 47.6 73.5 82.8 2 11.9 203.9
TS2-Net [13] DSL ECCV’22 51.1 76.9 85.6 1 11.7 213.6 - - - - - - 47.4 74.1 82.4 2 12.9 203.9
TABLE [10] DSL AAAI’23 52.3 78.4 85.2 1 11.4 215.9 - - - - - - 49.1 75.6 82.9 2 14.8 207.6
UATVR [33] DSL ICCV’23 49.8 76.1 85.5 2 12.3 211.4 - - - - - - - - - - - -
RAP [85] DSL ACL’24 - - - - - - - - - - - - 47.1 74.1 82.4 - 13.9 203.6
TGIF Charades YouCook2
RIVRL [86] - TCSVT’22 12.2 26.8 35.4 30 - 74.4 - - - - - - - - - - - -
X-Pool [26] - CVPR’22 - - - - - - 16.1 35.2 44.9 14 67.2 96.2 - - - - - -
AME-Net [87] - TOMM’22 - - - - - - - - - - - - 7.6 21.5 32.8 28 - 61.9
EMCL-Net [15] - NeurIPS’22 13.2 27.3 34.8 36 327.5 75.3 16.4 36.3 47.3 13 69.5 100.0 11.8 28.8 39.8 19 121.6 80.4
RoME [88] - CoRR’22 - - - - - - - - - - - - 6.3 16.9 25.2 53 - 48.4
SwAMP [89] - AISTATS’23 - - - - - - - - - - - - 9.4 24.9 35.3 22 - 69.6
MuMUR [79] - Inf. Retr. J.’23 - - - - - - 16.6 37.5 50.0 10 52.7 104.1 - - - - - -
UCoFiA [82] - ICCV’23 13.0 27.1 34.7 35 323.8 74.8 16.4 37.2 47.7 12 74.6 101.3 12.4 30.4 41.0 17 123.5 83.8
LTME [90] - ICASSP’24 10.6 24.1 32.2 - - 66.9 - - - - - - - - - - - -
BiC-Net [84] - TOMM’24 - - - - - - - - - - - - 8.7 23.9 33.5 31 - 66.1
EMCL-Net [15] DSL NeurIPS’22 14.8 29.4 36.6 31 317.3 80.8 18.2 38.2 49.3 11 63.1 105.7 12.6 31.7 42.5 17 104.5 86.8
TABLE X: Comparison with SOTA across all the datasets. --: unreported results. our reproduced results. IS: Inference Strategy. The highest $RSum$ in each block is marked in bold.
baseline 52.1 77.3 85.9 1 12.9 215.3
+VT 52.2 77.3 85.7 1 10.4 215.2
+TT 52.0 77.6 86.0 1 10.4 215.6
+VT+TT 53.2 77.7 85.3 1 10.0 216.2
TABLE XI: Ablation on architecture design.
Foundation Models
VT TT $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
VL L2 52.0 77.5 84.9 1 10.4 214.4
VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
TABLE XII: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
Captions Blip2 PG 51.2 76.2 85.0 1 11.2 212.4
Captions Blip2 L3.1 51.6 76.7 85.5 1 10.9 213.8
Captions VL2 L3.1 50.4 75.9 84.7 1 11.5 211.0
Tags VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
TABLE XIII: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
Appendix C Additional Ablations
In this section, we ablate some additional design choices of MAC-VR. First, we ablate $K$ that controls the number of concepts/latent factors. We evaluate within the following range $K\in{4,8,16,32}$. In Tab. VIII, the performance improves reaching the best result with $K=8$ and then decreases. This shows that a small number of concepts limits the ability to leverage fine-grained information, whereas a larger value reduces the dimensionality of each concept limiting the discriminability of the concept itself.
In Eq. 5, we described our training loss and we introduced two weight parameters $\alpha_{1}$ and $\alpha_{2}$ that control the importance of the contrastive loss $L_{C}$ to align the modality concepts and the proposed Alignment Loss $L_{A}$ to align the modality with the auxiliary modality concepts. We consider different values of $\alpha_{1}\in\{0.5,1.0,2.0,5.0\}$ and $\alpha_{2}\in\{0.0,0.5,1.0,2.0,5.0,10.0\}$. We find that the best $RSum$ is when $\alpha_{1}=1.0$ and $\alpha_{2}=1.0$ with $RSum=216.2$. The performance drops when $\alpha_{1}>2.0$ or $\alpha_{2}>2.0$, as shown in Tab. VIII. These results show that both losses must contribute to the learning phase equally to better align visual and textual concepts and so improve the final retrieval performance.
Appendix D Differences between tags extracted from different foundations models
As explained in Sec. IV-D, we considered different foundation models to extract visual and textual tags. More precisely, we considered Video-LLaMA [38] and VideoLLaMA2 [37] to extract visual tags and Llama2 [40] and Llama3.1 [95] to extract textual tags. We used the same parameters and same prompts to extract the tags by using all the considered foundations models.
Tab. VIII shows the statistics of extracted tags for MSR-VTT-9k. The total number of unique extracted visual tags by using Video-LLaMA [38] is much smaller than the total number obtained when using VideoLLaMA2 [37]. The same conclusion is reached for the total number of unique extracted textual tags by Llama2 [40] and Llama3.1 [95] respectively. This shows the lesser ability of older foundation models to generate tags compared with more recent versions. In particular, this is more evident when we focused on the visual tags, where not only the total number of unique tags is smaller but also the average number of tags per pairs is the same, meaning that many tags are shared among pairs and so there are less unique tags able to distinguish all the video/caption pairs.
Figure 9: Comparison of extracted visual and textual tags on the MSR-VTT dataset when using different foundation models. Blue and orange colours highlight meaningful tags.
For example, given the captions a cartoon character runs around inside of a video game and its corresponding video, we can see that Video-LLaMA and Llama2 hallucinate some visual tags such as snowboarding, desert, skiing, bird, climbing—definitely irrelevant to what appear in the video—and textual tags such as video game, running, character, game character that are already words that appear in the caption, therefore they do not add any additional information to better retrieve the correct video. On the contrary VideoLLaMA2 and Llama3.1 tend to extract tags that add additional information to the video and text. See Fig. 9 for more examples on all the considered datasets.
Appendix E How to Generate Auxiliary Captions.
In Sec. IV-D, we ablate the use of auxiliary captions instead of using tags. We generate these additional captions by extracting them directly from the video and paraphrasing the original caption. We consider different approaches to extract captions from video and text.
Visual Captions. We consider two different approaches to generate new captions from a video: Blip2 [34] and VideoLLaMA2 [37]. Following the same approach proposed in [101], we generate new captions by extracting the middle frame of each video and use Blip2 to generate a new caption. We use the same parameters of VideoLLaMA2 to extract captions as we did to extract visual tags in MAC-VR, as described in Sec. IV-B. However, we used a general prompt to ask VideoLLaMA2 to generate new captions. {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
You are a conversational AI agent. You typically look at a video and generate a new caption for a video. Generate 10 new captions. Give me a bullet list as output.
Textual Captions. We consider the paraphraser PEGASUS [98] and Llama3.1 [95] to paraphrase the original caption. PEGASUS [98] is a standard Transformer-based encoder-decoder method pre-trained on a massive text corpora with a novel pre-training objective called Gap Sentence Generation (GSG). Instead of using traditional language modeling, PEGASUS removes important sentences from a document (gap-sentences) and asks the model to predict these missing sentences. After the pre-training stage, PEGASUS is fine-tuned on specific summarization datasets to improve its performance on downstream tasks. The model becomes highly effective at generating concise and accurate summaries by leveraging its pre-training knowledge. We extract new captions from a caption by Llama3.1 [95] by giving as input a general prompt as we did to extract tags:
{mdframed}
[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: You are a conversational AI agent. You typically paraphrase sentences by using different words but keeping the same meaning.
Given the following sentence: 1) {}
Generate 10 different sentences that are a paraphrased version of the original sentence. Give me a bullet list as output.
ASSISTANT:
We randomly pick an extracted visual and textual caption as auxiliary inputs in MAC-VR during training, noting a similar length of input with our ‘sentence of tags’. In inference, we always pick the first caption in the set of the extracted ones. Some examples of the extracted captions with all the considered methods are shown in Fig. 10.
@@ Appendix D Differences between tags extracted from different foundations models
Appendix F Additional t-SNE plot.
In Fig. 12, we show the t-SNE plot of MAC-VR on the MSR-VTT test set when using modality-specific tags with/without our Alignment Loss $\mathcal{L}_{A}$. It is clear that the use of $\mathcal{L}_{A}$ helps to better distinguish the different concepts and have better clusters in the t-SNE plot. In Fig. 13 and Fig. 14, we show the t-SNE plot of visual and textual concepts without auxiliary modality-specific tags and when using only visual tags (Fig. 13) and textual tags (Fig. 14) with our Alignment Loss $\mathcal{L}_{A}$. From this, both tags used individually help to better align the visual and textual concepts. In particular, the visual tags help to better align the visual and textual concepts compared to the textual tags. A possible explanation for this is that the tags extracted from videos share the same modality as captions, which facilitates better alignment between visual and textual concepts. We leave this conclusion as possible inspiration for future works in this field.
@@
Fig. 15 to 17 show the t-SNE plot of MAC-VR of visual and textual concepts with/without auxiliary modality-specific tags across the other datasets. We sample $1,000$ pairs for TGIF and YouCook2. We can see that the use of auxiliary modality-specific tags help to better distinguish the different concepts and have better clusters in the t-SNE plot.
Appendix G Qualitative Results
In Fig. 11, we show additional qualitative results across all the datasets. It is evident that tags add additional information that is not described in the text. For example, the video associated to the caption a guy is in the water with his surfboard has visual (i.e. water activity, ocean, waves, surfing, outdoor activity) and textual (i.e. surfing, water, ocean, riding) tags that add additional information and boost the rank position by $3$. Similarly, in the YouCook2 example, the video is associated to a general mash the beans, tags help to add additional information, indeed visual tags such as boiling, homemade soup, boiling liquid, thickening sauce and textual tags such as such as crushing action, bean processing, pureeing, ingredient mashing add semantic information that definitely help the model in the retrieval process. Rarely, the extracted tags were not semantically relevant with what shown in the video or described in the caption. For instance, some visual and textual tags of the video associated to the caption the man blows out the candles. the girl puts the cake down. the camera pans to a man blowing out the candles on his cake. man blows candles out a plate is placed on the table are marriage partnership, love, love commitment as visual tags and class as textual tags. In other cases, visual and textual tags are too general, even though they are relevant to the video and caption, and so they do not add any additional semantic information. For example the video associated to the caption place wasabi next to the tuna are food, cooking, preparation techniques as visual tags and ingredients, serving, serving food. Even though these tags are relevant to the video and caption, they are too general and do not add any additional semantic knowledge.
Figure 11: Additional qualitative results across datasets. Blue and orange colours highlight some meaningful tags.
(a) w/ Modality-Specific Tags, w/o $\mathcal{L}_{A}$.
(b) w/ Modality-Specific Tags, w/ $\mathcal{L}_{A}$.
Figure 12: t-SNE plot of visual and textual concepts on the MSR-VTT-9k with/without the Alignment Loss $\mathcal{L}_{A}$ with using both visual and textual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Visual Tags, w/ $\mathcal{L}_{A}$.
Figure 13: t-SNE plot of visual and textual concepts on the MSR-VTT-9k without using auxiliary modality-specific tags and with using only visual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Textual Tags, w/ $\mathcal{L}_{A}$.
Figure 14: t-SNE plot of visual and textual concepts on the MSR-VTT-9k without using auxiliary modality-specific tags and with using only textual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
@@ Appendix G Qualitative Results
Figure 15: t-SNE plot of visual and textual concepts on the DiDeMo with/without using auxiliary modality-specific tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
@@
Figure 16: t-SNE plot of visual and textual concepts on the TGIF with/without using auxiliary modality-specific tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.

### v3

@@ Appendix F Additional Ablations
Table 15: Ablation on number of concepts $K$.
$\alpha_{C}$ $\alpha_{A}$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
@@
Table 16: Ablation on loss parameters $\alpha_{C}$ and $\alpha_{A}$.
In our training loss, we introduced two weight parameters $\alpha_{C}$ and $\alpha_{A}$ that control the importance of the contrastive loss $L_{C}$ to align the modality concepts and the proposed Alignment Loss $L_{A}$ to align the modality with the auxiliary modality concepts. We consider different values of $\alpha_{C}\in\{0.5,1.0,2.0,5.0\}$ and $\alpha_{A}\in\{0.0,0.5,1.0,2.0,5.0,10.0\}$. We find that the best $RSum$ is when $\alpha_{C}=1.0$ and $\alpha_{A}=1.0$ with $RSum=216.2$. The performance drops when $\alpha_{C}>2.0$ or $\alpha_{A}>2.0$, as shown in Tab. 16. These results show that both losses must contribute to the learning phase equally to better align visual and textual concepts and so improve the final retrieval performance.
Appendix G Differences between tags extracted from different foundations models
As explained in Sec. 4.3, we considered different foundation models to extract visual and textual tags. More precisely, we considered Video-LLaMA [Zhang et al.(2023a)Zhang, Li, and Bing] and VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] to extract visual tags and Llama2 [Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al.] and Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] to extract textual tags. We used the same parameters and same prompts to extract the tags by using all the considered foundations models. Video-LLaMA and Llama2 tend to hallucinate tags more often as shown in Tab. 17 and Fig. 9. More precisely, Tab. 17 shows the statistics of extracted tags for MSR-VTT-9k. The total number of unique extracted visual tags by using Video-LLaMA [Zhang et al.(2023a)Zhang, Li, and Bing] is much smaller than the total number obtained when using VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.]. The same conclusion is reached for the total number of unique extracted textual tags by Llama2 [Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al.] and Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] respectively. This shows the lesser ability of older foundation models to generate tags compared with more recent versions. In particular, this is more evident when we focused on the visual tags, where not only the total number of unique tags is smaller but also the average number of tags per pairs is the same, meaning that many tags are shared among pairs and so there are less unique tags able to distinguish all the video/caption pairs.
Figure 9: Comparison of extracted visual and textual tags on the MSR-VTT dataset when using different foundation models.
@@ Appendix G Differences between tags extracted from different foundations models
Video-LLaMA [Zhang et al.(2023a)Zhang, Li, and Bing] Visual 8,049 3,500 27,11 27,12
VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] Visual 63,383 12,118 27.69 27.83
Llama2 [Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al.] Textual 162,571 5,058 14.96 15.20
Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] Textual 320,351 8,326 27.17 26.52
@@
Table 17: Comparison of statistics of tags when using different foundation models on MSR-VTT-9k.
For example, given the captions a cartoon character runs around inside of a video game and its corresponding video, we can see that Video-LLaMA and Llama2 hallucinate some visual tags such as snowboarding, desert, skiing, bird, climbing—definitely irrelevant to what appear in the video—and textual tags such as video game, running, character, game character that are already words that appear in the caption, therefore they do not add any additional information to better retrieve the correct video. On the contrary VideoLLaMA2 and Llama3.1 tend to extract tags that add additional information to the video and text.
See Fig. 9 for more examples on all the considered datasets.
Appendix H How to Generate Auxiliary Captions.
In Sec. 4.3, we ablate the use of auxiliary captions instead of using tags. We generate these additional captions by extracting them directly from the video and paraphrasing the original caption. We consider different approaches to extract captions from video and text.
Visual Captions. We consider two different approaches to generate new captions from a video: Blip2 [Li et al.(2023b)Li, Li, Savarese, and Hoi] and VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.]. Following the same approach proposed in [Wang et al.(2024b)Wang, Li, Li, Yu, He, Chen, Pei, Zheng, Wang, Shi, Jiang, Li, Xu, Zhang, Huang, Qiao, Wang, and Wang], we generate new captions by extracting the middle frame of each video and use Blip2 to generate a new caption. We use the same parameters of VideoLLaMA2 to extract captions as we did to extract visual tags in MAC-VR. However, we used a general prompt to ask VideoLLaMA2 to generate new captions. You are a conversational AI agent. You typically look at a video and generate a new caption for a video. Generate 10 new captions. Give me a bullet list as output. Textual Captions. We consider the paraphraser PEGASUS [Zhang et al.(2020)Zhang, Zhao, Saleh, and Liu] and Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] to paraphrase the original caption. PEGASUS [Zhang et al.(2020)Zhang, Zhao, Saleh, and Liu] is a standard Transformer-based encoder-decoder method pre-trained on a massive text corpora with a novel pre-training objective called Gap Sentence Generation (GSG). Instead of using traditional language modeling, PEGASUS removes important sentences from a document (gap-sentences) and asks the model to predict these missing sentences. After the pre-training stage, PEGASUS is fine-tuned on specific summarization datasets to improve its performance on downstream tasks. The model becomes highly effective at generating concise and accurate summaries by leveraging its pre-training knowledge.
@@
We extract new captions from a caption by Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] by giving as input a general prompt as we did to extract tags: A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: You are a conversational AI agent. You typically paraphrase sentences by using different words but keeping the same meaning. Given the following sentence: 1) {} Generate 10 different sentences that are a paraphrased version of the original sentence. Give me a bullet list as output. ASSISTANT: We randomly pick an extracted visual and textual caption as auxiliary inputs in MAC-VR during training, noting a similar length of input to our ‘sentence of tags’. In inference, we always pick the first caption in the set of the extracted ones. Some examples of the extracted captions with all the considered methods are shown in Fig. 10.
Appendix I Additional t-SNE plot.
@@
Fig. 11 to 14 show the t-SNE plot of MAC-VR of visual and textual concepts with/without auxiliary modality-specific tags across all the other datasets. We sample $1,000$ pairs for Charades, TGIF and YouCook2. We can see that, also for these datasets, the use of auxiliary modality-specific tags helps to better distinguish the different concepts and in the t-SNE plot.
In Fig. 15, we show the t-SNE plot of MAC-VR on the MSR-VTT-9k test set when using modality-specific tags with/without our Alignment Loss $\mathcal{L}_{A}$. It is clear that the use of $\mathcal{L}_{A}$ helps to better distinguish the different concepts and have better clusters in the t-SNE plot. In Fig. 16 and Fig. 17, we show the t-SNE plot of visual and textual concepts without auxiliary modality-specific tags and when using only visual tags (Fig. 16) and textual tags (Fig. 17) with our Alignment Loss $\mathcal{L}_{A}$. From this, both tags used individually help to better align the visual and textual concepts. In particular, the visual tags help to better align the visual and textual concepts compared to the textual tags. A possible explanation for this is that the tags extracted from videos share the same modality as captions, which facilitates better alignment between visual and textual concepts. We leave this conclusion as possible inspiration for future works in this field.
@@
Figure 11: t-SNE plot of visual and textual concepts on DiDeMo without (left) and with using auxiliary modality-specific tags (right).
@@
Figure 12: t-SNE plot of visual and textual concepts on Charades without (left) and with using auxiliary modality-specific tags (right).